What is the value of x in the quadratic equation "4a^2=2a"
Answer by Alan3354(69443)

What is the value of x in the quadratic equation "4a^2=2a"
=====================
There's no x.
------------------
4a^2=2a
2a^2 = a
2a^2 - a = 0
a*(2a - 1) = 0
a = 0, a = 1/2